Please help with calculation
Homework 1b problems, due with Homework1 1) A persons metabolic rate is the rate at which the body consumes energy. Metaboli is important in studies of weight gain, dieting, and exercise. Below are the metabolic rates of 12 women who took part in a study of dieting. Metabolic rates are kilocalories (Cal) for a 24- hour period. c rate 913 995 1052 1124 1189 1204 1256 1347 1396 1418 1425 1502 A) By hand calculate the 5 number summary (show work) B) Now, please calculate the standard deviation (by hand) for a smaller set of women below 913. 995 1052 1124 1189 1204
Answer #1

a)below is 5 number summary:

minimum =913

quartile 1 =npth value =12*0.25 th value =4th value =(1124+1189)/2=1156.5

quartile 2 =npth value =12*0.5 th value =6th value =(1204+1256)/2=1230

quartile 3 =npth value =12*0.75 th value =9th value =(1396+1418)/2=1407

maximum=1502

B)

(x-i) -166.500 84.500 -27.500 S.No (X-x)2 913.000 995.000 1052.000 756.2500 1980.2500 11990.2500 15500.2500 65089.5000 13017.

from above std deviation=114.096